Blue

This package facilitates placing boxes at absolute positions on the L^{A}TEX page. There are several reasons why this might be useful, but the main one (or at least my motivating one) is to help produce a large-format conference poster. This package provides a single environment, which contains the text (or graphics, or table, or whatever) which is to be placed on the page, and which specifies where it is to be placed.

Blue

This package facilitates placing boxes at absolute positions on the L^{A}TEX page. There are several reasons why this might be useful, but the main one (or at least my motivating one) is to help produce a large-format conference poster. This package provides a single environment, which contains the text which is to be placed on the page, and which specifies where it is to be placed.

Blue

This package facilitates placing boxes at absolute positions on the L^{A}TEX page. There are several reasons why this might be useful, but the main one (or at least my motivating one) is to help produce a large-format conference poster. This package provides a single environment, which contains the text (or graphics, or table, or whatever) which is to be placed on the page, and which specifies where it is to be placed.